cost accounting

A Corporation sells its product for $17 per unit.  Its variable cost is $10 per unit, and total fixed costs are $800.  Assuming next period’s estimated sales are 300, calculate the following amounts:

a.   Degree of operating leverage

b.   Margin of safety in units

c.   Margin of safety in revenues
Answer #1

·         Contribution Margin per unit = Sales Price – Variable Cost per Unit = 17- 10 = 7 unit

·         Contribution Margin Ratio = Contribution Margin / Sales price = (7/17) * 100= 41.2%

·         Break-even in units= fixed costs/ Contribution Margin per unit=800/7=114.3 unit

·         Break-even Point in Dollars = fixed costs/ Contribution Margin Ratio= 800/41.2%=1941

·         Break even in units= 300+800/7=157 unit

·         Break even in dollars = 300+800/41.2%=2670

·         Sales = 114.3 units * 17 sales price per unit = 1938+300=2243

 

c-  Margin safety in dollars =sales-Break even sales = 2243-1941=302

 

·        Variable costs = 10 per unit * 114.3 units = 1143

 

a- Degree of operating leverage=(sales-variable costs) /(sales – variable costs-fixed costs)=(2243-114) / (2243-1143-800) = 3.7

 

b-  Margin of safety in units= Current Sales Units – Breakeven Point=157-114=43 unit


answered by: nonolll22